# Bargaining via Weber’s law

### v1

@@ 1 Introduction
We develop a bargaining theory based on Weber’s law (WL) [10, 11]: only those absolute differences $\Delta S$ between two stimuli are discriminated for which
$|\Delta S|\geq kS,\quad 1>k_{1}>0,\quad 1>k_{2}>0.$ (1)
where $S$ is the intensity of the baseline stimulus, and where $k>0$ is a constant that does not depend on $\Delta S$ and $S$ but depends on the nature of the stimulus and the perceiving individual. The minimal absolute difference $|\Delta S|=kS$ defines the concept of the just-noticeable-difference. Limitations of WL are discussed in [10, 12]. Eq. (1) does not depend on the units chosen to measure $S$. This is crucial for applications of (1) to bargaining.
WL relates to the Weber-Fechner law, which involves additional assumptions, but can be easier to verify in practice [13]. WL has traditionally been applied to the perception of weight, loudness, and brightness [10, 11]. More recently, the law was observed and validated on cognitive systems, e.g. the perception of temporal durations [14], or representation of numbers and numerosity in humans and animals [15, 16].
@@ 2 Bargaining set-up
As a proposer, each player acts as follows. Our first axiom is that the proposal is made on the Pareto line (3). The second axiom is that $\mathcal{P}_{1}$  and $\mathcal{P}_{2}$  hold (1) for relevant utility changes with constants $k_{1}$ and $k_{2}$, respectively. The utility $u_{1}$ of $\mathcal{P}_{1}$  is maximized provided that taking $u_{1}-d_{1}$ as the baseline in (1), and increasing it by $w_{1}$ is still relevant for $\mathcal{P}_{1}$  ($\frac{w_{1}}{u_{1}-d_{1}}\geq k_{1}$), and is acceptable for $\mathcal{P}_{2}$. The resulting utility $u_{1}^{*}$ reads:

### v2

@@ 1 Introduction
We develop a bargaining theory based on Weber’s law (WL) [10, 11]: only those absolute differences $\Delta S=S^{\prime}-S$ between two stimuli $S^{\prime}$ and $S$ are discriminated for which
$|\Delta S|\geq kS,\quad 1>k>0.$ (1)
where $S$ is the intensity of the baseline stimulus, and where $k>0$ is a constant that does not depend on $\Delta S$ and $S$ but depends on the nature of the stimulus and the perceiving individual. Baseline means that the stimulus $S$ came first, and is considered certain, i.e. $S^{\prime}$ is to be discriminated from $S$, and on the background of $S$. The minimal absolute difference $|\Delta S|=kS$ defines the concept of the just-noticeable-difference. Limitations of WL are discussed in [10, 12]. Eq. (1) does not depend on the units chosen to measure $S$. This is crucial for applications of (1) to bargaining.
WL relates to the Weber-Fechner law, which involves additional assumptions, but can be easier to verify in practice [13]. WL has traditionally been applied to the perception of weight, loudness, brightness, and time [10, 11, time_weber]. More recently, the law was observed and validated on cognitive systems, e.g. the perception of temporal durations [14], or representation of numbers and numerosity in humans and animals [15, 16].
@@ 2 Bargaining set-up
As a proposer, each player acts as follows. Our first axiom is that the proposal is made on the Pareto line (3). The second axiom is that $\mathcal{P}_{1}$  and $\mathcal{P}_{2}$  hold (1) for relevant utility changes with constants $k_{1}$ and $k_{2}$, respectively. This means that the utility $u_{1}$ of $\mathcal{P}_{1}$  is maximized in two steps. First $u_{1}-d_{1}$ is taken as the baseline in (1), so that it is still possible to increase it by $w_{1}$. This increase is relevant for $\mathcal{P}_{1}$  (i.e. $\frac{w_{1}}{u_{1}-d_{1}}\geq k_{1}$ according to (1)), and is acceptable for $\mathcal{P}_{2}$(i.e. $\frac{h(u_{1})-h(u_{1}+w_{1})}{h(u_{1})-d_{2}}\leq k_{2}$). The resulting maximal utility $u_{1}^{*}$ is the penultimate noticeable step before hitting the acceptance level of the opponent: